# A Systematic Survey on Large Language Models for Evolutionary Optimization: From Modeling to Solving

### v2

@@ V-B3 Algorithm Configuration
The performance of EAs is strongly influenced by hyperparameter and operator configurations. Traditionally, configuration has been studied through static tuning [183] and dynamic control [184]. In recent years, dynamic configuration has become particularly prominent, largely driven by RL methods [185, 186, 187], given the natural alignment between the Markov Decision Process (MDP) framework [188] and adaptive configuration. However, RL-based approaches usually require problem-specific training, leading to high computational costs and limited generalization. Leveraging their reasoning abilities, LLMs provide a promising alternative by analyzing optimization features and recommending adaptive configurations without extensive retraining.
Several representative works illustrate this direction. Kramer et al. introduced an LLM-based feedback loop for both static tuning [135] and dynamic control [142] of evolutionary strategies. Custode et al. [139] applied an OPRO-based mechanism, feeding optimization trajectories back to the LLM for step-size control, and additionally prompted the model to justify its parameter updates. Zhang et al. [145] addressed the redundancy of trajectory-based feedback by proposing a meta-prompt built on features from the solution space, decision space, and optimization process. They further integrated traditional adaptive control and prior optimization experience to enable LLM-guided adaptive operator selection.
It is worth noting that some algorithm generation frameworks also involve hyperparameter tuning. For example, EoH [154] employs prompt templates for parameter tuning during heuristic mutation, while LLaMEA [157] integrates hyperparameter optimization to reduce iteration costs. However, the primary contribution of these works lies in algorithm generation rather than EA configuration, and thus they fall outside the strict scope of configuration research.
@@ V-B4 Evaluation
In computationally expensive optimization problems, surrogate models are widely used to approximate costly real-world evaluations and alleviate computational bottlenecks. Recently, LLMs have been explored as powerful tools for surrogate modeling, leveraging their pattern recognition and reasoning capabilities to improve evaluation efficiency.
A pioneering study by Hao et al. [136] introduced an LLM-based surrogate that reframes model-assisted selection as classification and regression, enabling evaluation of new solutions from historical data without additional training. Building on this idea, Nguyen et al. [137] developed the LICO framework to address data scarcity in domains such as molecular optimization, where direct LLM application is challenging. Beyond acting as surrogates, LLMs have also been employed for managing surrogate-assisted algorithms. Rios et al. [138] incorporated LLMs into the selection and training of surrogate models for engineering optimization. Xie et al. [149] further proposed LLM-SAEA, a framework that uses collaborative experts to dynamically select surrogate models and infill criteria, reducing reliance on extensive domain expertise.
More recently, Zhang et al. [147] extended this line of research to multi-task optimization, treating LLMs as meta-surrogates capable of facilitating knowledge transfer across tasks through unified token sequence representations. These studies collectively highlight the potential of LLMs not only as direct surrogates but also as adaptive coordinators in surrogate-assisted optimization, though further work is required to assess their robustness and scalability in large-scale applications.
@@ V-C1 Algorithm Selection
The goal of algorithm selection is to identify the most suitable algorithm from a portfolio for each problem instance. This problem arises from the significant performance disparities among algorithms, which often exhibit different optimization behaviors and application domains [7]. Traditionally, algorithm selection is formulated as a machine learning task, where a selector is trained on problem instances and their features to predict the best algorithm for unseen cases [189]. The process typically involves two stages: feature extraction and selector construction. The quality of extracted features directly affects the quality of training samples and thus the performance of the selector. Commonly used features include basic statistical descriptors of the problem [190] as well as optimization characteristics obtained from algorithm runs [191]. Selector construction can then be approached through classification [192], regression [193], or hybrid models [194], offering a wide range of technical solutions for this task.
@@
Based on this workflow, recent studies have applied LLMs to both feature extraction and selector construction. Wu et al. introduced the AS-LLM framework [153], the first to bring LLMs into algorithm selection. AS-LLM leverages the code understanding capabilities of LLMs to automatically extract high-dimensional algorithm features from code or descriptive text. Its core representation module combines LLM-extracted features with a feature selection mechanism to identify the most relevant subset. The resulting algorithm features and problem features are then processed in separate networks to avoid premature interaction, and a similarity calculation module determines the final choice by evaluating the match between problems and algorithms. In contrast, Zhang et al. proposed InstSpecHH [176], which emphasizes selector construction. After a Euclidean distance–based pre-selection step narrows the candidate algorithms, the natural language description of the problem instance and the pre-selected algorithms are fed into the LLM. This approach exploits the LLM’s semantic reasoning ability to achieve more context-aware selection. It is worth noting that some related works involve solver selection [71, 79], where LLMs are integrated into broader workflows such as optimization modeling or end-to-end pipeline design. While relevant, these works differ from the classical definition of algorithm selection and are generally not categorized as such.
@@ V-C2 Algorithm Generation
In contrast to algorithm configuration and selection, algorithm generation aims to automatically design optimization algorithms from problem characteristics without relying on pre-defined component or algorithm pools. Reinforcement learning–based approaches have made progress in this direction, but they still depend on manually specified component or symbol sets [35]. For instance, the GSF framework [195] can generate algorithms within a generic EA template, yet each component pool must be pre-specified. Chen et al. proposed SYMBOL [17], which removed the component pool constraint but continued to rely on a manually defined symbol set. LLMs, with their broad knowledge bases and strong generative capabilities, offer a promising way to overcome these limitations. By directly producing algorithmic structures and logic, LLMs bypass the need for pre-specified pools, making algorithm generation one of the most active branches of LLM-assisted optimization. Research in this area has evolved from early single-step generation toward more advanced iterative generation frameworks [153].
@@
Single-step generation represents the initial exploration of this field. Pluhacek et al. [150] used GPT-4 to decompose and recombine six metaheuristic algorithms, creating new hybrid variants. Zhong et al. [196] combined GPT-3.5 with a prompt engineering framework to generate ZSO, a novel metaheuristic, in a single step. These studies demonstrated the feasibility of LLM-driven algorithm generation but also revealed the limitations of single-step methods, which depend heavily on the model’s inherent capabilities. Consequently, current research has shifted toward iterative generation approaches, which provide greater adaptability and robustness.
@@
Iterative generation has since become the dominant paradigm, with research progressing rapidly from foundational studies to real-world applications. A pioneering work in this direction is Funsearch [151], which integrates LLMs with evolutionary algorithms to iteratively generate program snippets in function space. Building on this idea, Liu et al. proposed AEL [152] and EoH [154], where heuristic concepts co-evolve with code, achieving superior performance on combinatorial optimization benchmarks. To support this line of work, they developed the LLM4AD platform [167] and analyzed the multimodal and rugged fitness landscape of LLM-assisted algorithm search [174]. More recently, they fine-tuned LLMs specifically for algorithm design, introducing diversity-aware ranking, sampling strategies, and direct preference optimization, demonstrating that fine-tuned models can significantly outperform generic ones [178]. The EoH framework has since been extended into MEoH [163] for multi-objective heuristic search and EoH-S [179] for co-evolving complementary sets of algorithms, with applications in edge server task scheduling [162], Bayesian optimization [161], and adversarial attack design [158].
@@
In parallel, Van Stein and Bäck developed the LLaMEA series [157, 197], which integrates LLMs into evolutionary algorithms for tasks such as mutation and selection, enabling the generation of advanced heuristics. To better characterize the generation process, they introduced code evolution graphs [169] and behavior space analysis [177], and built the BLADE framework [171] to provide standardized evaluation for LLM-driven algorithm generation. Extensions such as LLaMEA-HPO [164] incorporated hyperparameter optimization to improve efficiency, while applications have spanned Bayesian optimization [175] and photonic structure design [170].
@@
Beyond these two main lines, several studies have explored novel search paradigms. Ye et al. proposed ReEvo [155], which treats the LLM as a hyper-heuristic and introduces a reflective evolution mechanism. Zheng et al. developed MCTS-AHD [168], employing Monte Carlo Tree Search to explore heuristic space more thoroughly and avoid local optima. Pham et al. proposed HSEvo [166], combining harmony search with genetic algorithms and introducing diversity metrics to balance exploration and exploitation. Together, these efforts highlight the shift from single-step demonstrations toward robust iterative frameworks that can adaptively generate increasingly sophisticated optimization algorithms.
@@ V-D Challenges
When applying LLMs to optimization solving, methods at different levels exhibit distinct characteristics. The direct use of LLMs as optimizers, exemplified by OPRO [28], marked the initial integration of LLMs with optimization. Low-level approaches, such as LMEA [127] and LMX [125], have been deployed in domains like NAS and molecular discovery. High-level methods for algorithm generation, including EoH [154] and LLaMEA [157], currently represent the most active research frontier. While these studies have advanced the synergy between LLMs and optimization, several bottlenecks remain:
  • LLMs as Optimizers: Direct, trajectory-based optimization with LLMs often struggles to outperform classical algorithms [29] and performs poorly with smaller models that lack sufficient reasoning ability [117]. Moreover, trajectory-based iteration is difficult to extend to population-based optimization, since the explosion of trajectory data easily exceeds the LLM’s context capacity. Future work should explore alternative iterative forms that replace raw trajectories with more structured representations of optimization history.
  • Low-level LLMs for Optimization Algorithms: While the use of LLMs as evolutionary operators has shown promise in engineering, biology, and chemistry, research on initialization and algorithm configuration remains nascent. LLMs are difficult to apply to data-intensive tasks like large-scale initialization [143], and they often display conservative behavior in parameter control [139], failing to provide sufficient exploration. The underlying challenge is the vastness of the decision space. Future directions should focus on reducing decision-making pressure while preserving accuracy.
  • High-level LLMs for Optimization Algorithms: Major issues arise in both computational cost and innovation. Frameworks such as EoH [154] and LLaMEA [157] treat algorithms as individuals in an EA, which requires frequent LLM calls and costly evaluations during evolution. In terms of innovation, current methods mainly recombine existing strategies and remain limited in generating fundamentally novel approaches. Achieving scalable, low-cost search and fostering true algorithmic creativity are key open challenges for this line of research.
@@ VI-A Computer Science
In computer science, LLM-assisted optimization has been explored in multiple areas, with neural architecture search [198] and computer security [158] standing out as representative examples. Other emerging applications [199, 200, 201, 202] further illustrate the breadth of this direction.
NAS is a core problem in machine learning, where EAs have been widely applied [203]. Integrating LLMs offers new opportunities. While a few studies have attempted to treat LLMs as direct optimizers [204], most research emphasizes low-level assistance. Chen et al. [198] pioneered the use of LLMs as crossover and mutation operators within EAs, where crossover generates new architectures through few-shot prompting and mutation introduces diversity via mixed-temperature sampling. Building on this, Nasir [205] applied the approach to Quality-Diversity algorithms, achieving more efficient exploration with a dual-archive mechanism. Morris et al. [133] enhanced search diversity using role-based prompts, while Mo et al. [206] extended the paradigm to graph neural networks. In addition, LLMs have been explored for initialization: Yu et al. [126] leveraged pre-training, and Jawahar et al. [128] employed few-shot learning to provide high-quality initial suggestions.
In security-related optimization, LLMs have been used both as standalone optimizers and as auxiliary operators. Jiang et al. [119] applied LLMs directly to iteratively generate jailbreak attack suffixes with self-reflection. Complementary work explored low-level assistance [207, 208], where LLMs act as heuristic operators to guide jailbreak prompt optimization, improving both efficiency and transferability. Guo et al. [158] investigated LLM-enabled optimization of the generation function responsible for producing adversarial samples, supporting continuous innovation and evaluation of attack strategies.
Beyond NAS and security, LLM-based optimization has also been applied to emerging tasks. Wang et al. [199] treated data augmentation as an optimization problem, representing augmentation strategies as editable natural language instructions. By simulating crossover and mutation, LLMs intelligently explored strategies tailored to long-tailed distributions. Li et al. [200] focused on GPU computing, where LLMs iteratively generated and optimized CUDA kernel code, achieving systematic performance improvement without human intervention.

### v3

@@ V-B3 Algorithm Configuration
The performance of EAs is strongly influenced by hyperparameter and operator configurations. Traditionally, configuration has been studied through static tuning [huang2019survey] and dynamic control [karafotias2014parameter]. In recent years, dynamic configuration has become particularly prominent, largely driven by RL methods [chen2020self, pei2024learning, zhang2024improved], given the natural alignment between the Markov Decision Process (MDP) framework [puterman1990markov] and adaptive configuration. However, RL-based approaches usually require problem-specific training, leading to high computational costs and limited generalization. Leveraging their reasoning abilities, LLMs provide a promising alternative by analyzing optimization features and recommending adaptive configurations without extensive retraining.
Several representative works illustrate this direction. Kramer et al. introduced an LLM-based feedback loop for both static tuning [kramer2024large] and dynamic control [kramer2024llama] of evolutionary strategies. Custode et al. [custode2024investigation] applied an OPRO-based mechanism, feeding optimization trajectories back to the LLM for step-size control, and additionally prompted the model to justify its parameter updates. Zhang et al. [zhang2025laos] addressed the redundancy of trajectory-based feedback by proposing a meta-prompt built on features from the solution space, decision space, and optimization process. They further integrated traditional adaptive control and prior optimization experience to enable LLM-guided adaptive operator selection.
It is worth noting that some algorithm generation frameworks also involve hyperparameter tuning. For example, EoH [liu2024evolution] employs prompt templates for parameter tuning during heuristic mutation, while LLaMEA [van2024llamea] integrates hyperparameter optimization to reduce iteration costs. However, the primary contribution of these works lies in algorithm generation rather than EA configuration, and thus they fall outside the strict scope of configuration research.
@@ V-B4 Evaluation
In computationally expensive optimization problems, Surrogate Models (SMs) [liu2023surrogate] are widely used to approximate costly real-world evaluations and alleviate computational bottlenecks. Recently, LLMs have been explored as powerful tools for surrogate modeling, leveraging their pattern recognition and reasoning capabilities to improve evaluation efficiency.
A pioneering study by Hao et al. [hao2024large] introduced an LLM-based surrogate that reframes model-assisted selection as classification and regression, enabling evaluation of new solutions from historical data without additional training. Building on this idea, Nguyen et al. [nguyen2024lico] developed the LICO framework to address data scarcity in domains such as molecular optimization, where direct LLM application is challenging. Beyond acting as surrogates, LLMs have also been employed for managing surrogate-assisted algorithms. Rios et al. [rios2024large] incorporated LLMs into the selection and training of SMs for engineering optimization. Xie et al. [xie2025large] further proposed LLM-SAEA, a framework that uses collaborative experts to dynamically select SMs and infill criteria, reducing reliance on extensive domain expertise.
More recently, Zhang et al. [zhang2025large] extended this line of research to multi-task optimization, treating LLMs as meta-surrogates capable of facilitating knowledge transfer across tasks through unified token sequence representations. These studies collectively highlight the potential of LLMs not only as direct surrogates but also as adaptive coordinators in surrogate-assisted optimization, though further work is required to assess their robustness and scalability in large-scale applications.
@@ V-C1 Algorithm Selection
The goal of algorithm selection is to identify the most suitable algorithm from a portfolio for each problem instance. This problem arises from the significant performance disparities among algorithms, which often exhibit different optimization behaviors and application domains [kerschke2019automated]. Traditionally, algorithm selection is formulated as a machine learning task, where a selector is trained on problem instances and their features to predict the best algorithm for unseen cases [wu2025towards]. The process typically involves two stages: feature extraction and selector construction. The quality of extracted features directly affects the quality of training samples and thus the performance of the selector. Commonly used features include basic statistical descriptors of the problem [hutter2014algorithm] as well as optimization characteristics obtained from algorithm runs [nudelman2004understanding]. Selector construction can then be approached through classification [abdulrahman2018speeding], regression [xu2008satzilla], or hybrid models [fehring2022harris], offering a wide range of technical solutions for this task.
@@
Based on this workflow, recent studies have applied LLMs to both feature extraction and selector construction. Wu et al. introduced the AS-LLM framework [wu2023large], the first to bring LLMs into algorithm selection. AS-LLM leverages the code understanding capabilities of LLMs to automatically extract high-dimensional algorithm features from code or descriptive text. Its core representation module combines LLM-extracted features with a feature selection mechanism to identify the most relevant subset. The resulting algorithm features and problem features are then processed in separate networks to avoid premature interaction, and a similarity calculation module determines the final choice by evaluating the match between problems and algorithms. In contrast, Zhang et al. proposed InstSpecHH [zhang2025llm], which emphasizes selector construction. After a Euclidean distance–based pre-selection step narrows the candidate algorithms, the natural language description of the problem instance and the pre-selected algorithms are fed into the LLM. This approach exploits the LLM’s semantic reasoning ability to achieve more context-aware selection. It is worth noting that some related works involve solver selection [ahmaditeshnizi2024optimus, ahmaditeshnizi2024optimus03], where LLMs are integrated into broader workflows such as optimization modeling or end-to-end pipeline design. While relevant, these works differ from the classical definition of algorithm selection and are generally not categorized as such.
@@ V-C2 Algorithm Generation
In contrast to algorithm configuration and selection, algorithm generation aims to automatically design optimization algorithms from problem characteristics without relying on pre-defined component or algorithm pools. Reinforcement learning–based approaches have made progress in this direction, but they still depend on manually specified component or symbol sets [ma2025toward]. For instance, the GSF framework [yi2022automated] can generate algorithms within a generic EA template, yet each component pool must be pre-specified. Chen et al. proposed SYMBOL [chen2024symbol], which removed the component pool constraint but continued to rely on a manually defined symbol set. LLMs, with their broad knowledge bases and strong generative capabilities, offer a promising way to overcome these limitations. By directly producing algorithmic structures and logic, LLMs bypass the need for pre-specified pools, making algorithm generation one of the most active branches of LLM-assisted optimization. Research in this area has evolved from early single-step generation toward more advanced iterative generation frameworks [wu2023large].
@@
Single-step generation represents the initial exploration of this field. Pluhacek et al. [pluhacek2023leveraging] used GPT-4 to decompose and recombine six metaheuristic algorithms, creating new hybrid variants. Zhong et al. [zhong2024leveraging] combined GPT-3.5 with a prompt engineering framework to generate ZSO, a novel metaheuristic, in a single step. These studies demonstrated the feasibility of LLM-driven algorithm generation but also revealed the limitations of single-step methods, which depend heavily on the model’s inherent capabilities. Consequently, current research has shifted toward iterative generation approaches, which provide greater adaptability and robustness.
Iterative generation has since become the dominant paradigm, with research progressing rapidly from foundational studies to real-world applications. A pioneering work in this direction is Funsearch [romera2024mathematical], which integrates LLMs with evolutionary algorithms to iteratively generate program snippets in function space. Building on this idea, Liu et al. proposed AEL [liu2023algorithm] and EoH [liu2024evolution], fundamentally shifting the focus from purely code-based search to the co-evolution of abstract heuristic concepts and executable programs. EoH’s core innovation lies in its dual representation coevolutionary mechanism: it represents heuristic ideas as natural language thoughts and strategically leverages the LLM as a powerful translation engine to derive corresponding executable code snippets. The critical strength of EoH stems from this tight coupling and coevolution between high level concepts and low level implementations. This paradigm allows the evolutionary search to effectively explore the abstract algorithmic design space through the evolution of thoughts, which is significantly more efficient and creative than searching the pure code space, while simultaneously ensuring implementation fidelity. Consequently, EoH demonstrates superior performance and efficiency on challenging combinatorial optimization benchmarks, generating high performance heuristics that significantly outperform widely used handcrafted baseline algorithms.
@@
To support this line of work, they developed the LLM4AD platform [liu2024llm4ad] and analyzed the multimodal and rugged fitness landscape of LLM-assisted algorithm search [liu2025fitness]. More recently, they fine-tuned LLMs specifically for algorithm design, introducing diversity-aware ranking, sampling strategies, and direct preference optimization, demonstrating that fine-tuned models can significantly outperform generic ones [liu2025fine]. The EoH framework has since been extended into MEoH [yao2025multi] for multi-objective heuristic search and EoH-S [liu2025eoh] for co-evolving complementary sets of algorithms, with applications in edge server task scheduling [yatong2024ts], Bayesian optimization [yao2024evolve], and adversarial attack design [guo2024autoda].
@@
In parallel, Van Stein and Bäck developed the LLaMEA series [van2024llamea, van2025llamea], which represents a distinct trajectory by integrating LLMs directly as intelligent operators into the evolutionary algorithms for tasks such as mutation and selection, enabling the generation of advanced heuristics. In contrast to EoH’s focus on concept code coevolution, LLaMEA utilizes the LLM primarily to refine code within a standard evolutionary loop based on performance metrics and feedback from runtime evaluations. To better characterize the generation process, they introduced code evolution graphs [van2025code] and behavior space analysis [van2025behaviour], and built the BLADE framework [van2025blade] to provide standardized evaluation for LLM-driven algorithm generation. Extensions such as LLaMEA-HPO [van2024loop] incorporated hyperparameter optimization to improve efficiency, while applications have spanned Bayesian optimization [li2025llamea] and photonic structure design [yin2025optimizing].
@@
Beyond these two main lines, several studies have explored novel search paradigms [sun2024autosat, huang2025autonomous, zhang2024understanding, yin2025controlling, li2025llm, ling2025complex]. Ye et al. proposed ReEvo [ye2024reevo], which treats the LLM as a hyper-heuristic and introduces a reflective evolution mechanism. Zheng et al. developed MCTS-AHD [zheng2025monte], employing Monte Carlo Tree Search to explore heuristic space more thoroughly and avoid local optima. Pham et al. proposed HSEvo [dat2025hsevo], combining harmony search with genetic algorithms and introducing diversity metrics to balance exploration and exploitation. Together, these efforts highlight the shift from single-step demonstrations toward robust iterative frameworks that can adaptively generate increasingly sophisticated optimization algorithms.
@@ V-D Challenges
When applying LLMs to optimization solving, methods at different levels exhibit distinct characteristics. The direct use of LLMs as optimizers, exemplified by OPRO [yang2023large], marked the initial integration of LLMs with optimization. Low-level approaches, such as LMEA [liu2024large] and LMX [meyerson2024language], have been deployed in domains like NAS and molecular discovery. High-level methods for algorithm generation, including EoH [liu2024evolution] and LLaMEA [van2024llamea], currently represent the most active research frontier. While these studies have advanced the synergy between LLMs and optimization, several bottlenecks remain:
  • LLMs as Optimizers: Direct, trajectory-based optimization with LLMs often yields suboptimal performance compared to classical metaheuristics [huang2024exploring], and is ineffective with smaller models [zhang2024revisiting]. The core challenge is the conflict between iteration and LLM context efficiency. While context window capacity is no longer the bottleneck, non-structured, lengthy optimization trajectories severely degrade reasoning efficacy due to attention dispersion and the middle-loss effect. This context burden is exponentially exacerbated when scaling to population-based EC. Future work must therefore explore alternative iterative forms, replacing raw trajectories with structured and compressed optimization history representations to efficiently guide LLM decision-making.
  • Low-level LLMs for Optimization Algorithms: While the use of LLMs as evolutionary operators has shown promise in engineering, biology, and chemistry, research on initialization and algorithm configuration remains nascent. LLMs are difficult to apply to data-intensive tasks like large-scale initialization [zhao2025can], and they often display conservative behavior in parameter control [custode2024investigation], failing to provide sufficient exploration. The underlying challenge is the vastness of the decision space. Future directions should focus on reducing decision-making pressure while preserving accuracy.
  • High level LLMs for Optimization Algorithms: Major issues arise in both computational cost and innovation. Frameworks such as EoH [liu2024evolution] and LLaMEA [van2024llamea] treat algorithms as individuals in an EA, which requires frequent LLM calls and costly evaluations during evolution. In terms of innovation, current methods mainly recombine existing strategies and remain limited in generating fundamentally novel approaches. The root of this challenge lies in the difficulty of precise, hierarchical design: LLMs struggle to faithfully decompose a natural language problem statement and accurately map that deep understanding to fine grained design decisions, including the top level algorithmic framework, operator construction, and parameter settings. Achieving scalable, low cost search and fostering true algorithmic creativity are key open challenges for this line of research.
@@ VI-A Computer Science
In computer science, LLM-assisted optimization has been explored in multiple areas, with neural architecture search [chen2023evoprompting] and computer security [guo2024autoda] standing out as representative examples. Other emerging applications [wang2024llm, li2025cuda, ma2024llamoco, wang2025agenttts] further illustrate the breadth of this direction.
NAS is a core problem in machine learning, where EAs have been widely applied [liu2021survey]. Integrating LLMs offers new opportunities. While a few studies have attempted to treat LLMs as direct optimizers [zhong2024large], most research emphasizes low-level assistance. Chen et al. [chen2023evoprompting] pioneered the use of LLMs as crossover and mutation operators within EAs [yu2025llm], where crossover generates new architectures through few-shot prompting and mutation introduces diversity via mixed-temperature sampling. Building on this, Nasir [nasir2024llmatic] applied the approach to Quality-Diversity algorithms, achieving more efficient exploration with a dual-archive mechanism. Morris et al. [morris2024llm] enhanced search diversity using role-based prompts, while Mo et al. [mo2025autosgnn] extended the paradigm to graph neural networks. In addition, LLMs have been explored for initialization: Yu et al. [yu2023gpt] leveraged pre-training, and Jawahar et al. [jawahar2023llm] employed few-shot learning to provide high-quality initial suggestions.
In security-related optimization, LLMs have been used both as stand-alone optimizers and as auxiliary operators. Jiang et al. [jiang2024optimizable] applied LLMs directly to iteratively generate jailbreak attack suffixes with self-reflection. Complementary work explored low-level assistance [li2025efficient, yu2024llm], where LLMs act as heuristic operators to guide jailbreak prompt optimization, improving both efficiency and transferability. Guo et al. [guo2024autoda] investigated LLM-enabled optimization of the generation function responsible for producing adversarial samples, supporting continuous innovation and evaluation of attack strategies.
Beyond NAS and security, LLM-based optimization has also been applied to emerging tasks. Wang et al. [wang2024llm] treated data augmentation as an optimization problem, representing augmentation strategies as editable natural language instructions. By simulating crossover and mutation, LLMs intelligently explored strategies tailored to long-tailed distributions. Li et al. [li2025cuda] focused on GPU computing, where LLMs iteratively generated and optimized CUDA kernel code, achieving systematic performance improvement without human intervention.